OpenStudy (anonymous):

PLEASE CHECK MY ANSWERS? "/" MEANS THAT IT IS THE ANSWER I CHOSE. 1. Use the Distributive Property to find the price of 9 CDs that cost $13.99 each. A: $126.00 B: $117.00 C: $126.90 D: $125.91 /// 2. Teesha is in the French club. There are 39 students in the club. The French teacher will pick 5 students at random to guide visiting students from France. What is the probability that Teesha will not be picked as a guide? A: 5/39 B: 34/39 //// C: 5/44 D: 44/39 WILL BE CONTINUED...

OpenStudy (haichi):

2-A

OpenStudy (haichi):

1-D

OpenStudy (hunus):

Number two is B

OpenStudy (haichi):

how?

OpenStudy (hunus):

It's asking the probability that she won't be picked

OpenStudy (haichi):

o nvm

OpenStudy (haichi):

thnx

OpenStudy (haichi):

Im make mistakes some times sorry brandon :'(

OpenStudy (haichi):

I*

OpenStudy (anonymous):

3. You roll a blue number cube and a green number cube. Find P a number greater than 2 on the blue cube and an odd number on the green cube. A: 1/4 //// B: 1/6 C: 1/12 D: 1/3 4. You can fly from Seattle to Boise for $137.50, from Boise to Witchita for $92.50 and from Witchita to Austin for $162.50. Use mental math to find the total cost of the flight from Seattle to Austin. A: $402.50 B: $392.50 //// C: $394.50 D: $391.50 5. Name the property the equation illustrates. 0 + x = x A: Identity Property Of Addition //// B: Inverse Property Of Multiplication C: Multiplication Property Of 0 D: Commutative Property Of Addition 6. You roll a standard number cube. Find P(3 or 4). A: 1/3 ///// B: 1/2 C: 5/6 D: 1/6 7. You roll a standard number cube. Find P(number greater than 2). A: 1/3 B: 1 C: 2/3 //// D: 5/6 8. You roll a pair of number cubes. What is the probability of getting even numbers on both cubes? A: 1/2 //// B: 1/4 C: 1 D: 1/6 9. Simplify the expression. (–2 – c)(–1). A: -2 + c B: 2 - c C: 2 + c ///// D: -2 -c 10. Simplify 7(499) using the Distributive Property. A: 3514 B: 3493 ///// C: 3500 D: 3486 There will be 10 more.

OpenStudy (anonymous):

@Haichi , it's okay :D

OpenStudy (anonymous):

This is number nine...Because all I see are weird diamond symbols for that problem. Here is the question and the answer that I got.

OpenStudy (haichi):

5 is C

OpenStudy (haichi):

4 just add them all up

OpenStudy (haichi):

6 is A

OpenStudy (haichi):

7=C

OpenStudy (haichi):

8=A

OpenStudy (haichi):

10=7*499

OpenStudy (anonymous):

For #4 -- I did add them all up. The answer $392.50 is what I got.

OpenStudy (anonymous):

But thank you for the others :)

OpenStudy (haichi):

np

OpenStudy (anonymous):

11. You select a card at random from the cards that make up the word “replacement”. Without replacing the card, you choose a second card. Find the probability of choosing a vowel and then not a vowel. There is 1 letter for each card. A: 14/55 B: 28/121 //// C: 1/11 D: 4/121 12. You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($5 then $10). A: 1/9 B: 1/72 C: 2/27 D: 1/12 ///// 13. Simplify the expression. (–5 – c)(–1). A: -5 + c B: -5 - c C: 5 + c ////// D: 5 - c 14. SEE PHOTO ATTACHED. 15. Name the property the equation illustrates. 9 + (–9) = 0. A: Inverse Property Of Multiplication B: Identity Property Of Addition C: Addition Property Of 0 D: Inverse Property Of Addition ///// 16. You roll a standard number cube. Find P(number greater than 1). A: 1/6 B: 6/5 C: 5/6 ///// D: 1 17. The probability that a new spark plug is defective is 0.06. You need two new spark plugs. What is the probability that both spark plugs you buy are defective? A: 0.0036 ///// B; 0.06 C: 0.0030 D: 0.12 18. A cell phone company orders 500 new phones from a manufacturer. If the probability of a phone being defective is 2.6%, predict how many of the phones are likely to be defective. Round to the nearest whole number. A: 16 phones B; 130 phones C: 11 phones D: 13 phones //// 19. At the grocery store, you buy a carton of milk for $1.25, a loaf of bread on sale for $0.35, and a box of cereal for $3.75. Use mental math to find the total cost of the groceries. A: $4.75 B: $5.75 C: $5.35 ///// D: $5.85 20. What is the total cost if you buy 3 goldfish for $1.90 each, 3 angelfish for $6.10 each, and 12 neon tetras for $1.53 each? A: $66.36 B: $60.36 C: $39.36 D: $42.36 //////

OpenStudy (anonymous):

@Haichi Here is the rest. If there are any problems you can't read because weird symbols are there, let me know.

OpenStudy (haichi):

ok

OpenStudy (triciaal):

this is too long for number 3 I think it should be D 1/3 when you "and" the conditions you multiply the Probability >2 = 4/6 = 2/3 odd = 3/6 = 1/2 2/3*1/2 = 1/3

OpenStudy (anonymous):

@Haichi have you confirmed the answers yet? :)

OpenStudy (shadowlegendx):

@brandonfrabel are you using @haichi as your slave? ._.

OpenStudy (anonymous):

I'm not using him as my slave, haha. I ask him if he wants to help and he does and I reward him with a medal. Although he's such a big amount of help that I should send him real money for it lolol xD

OpenStudy (shadowlegendx):

@Haichi Just remember that giving out plain answers is against the CoC. You need to explain the answer and the question asker(That's you @brandonfrabel) must participate. If not, this is just giving out free answers with no learning involved.

OpenStudy (anonymous):

He isn't just giving me plain answers. I've already determined the answers and am simply asking Haichi to confirm whether or not the answers are correct. If I get a false answer, we'd elaborate and work out the problem together.

OpenStudy (anonymous):

@Hero can u tell me if my answers are correct

OpenStudy (anonymous):

We can all tell you if you're right or not, just show us what you did to get to the answer, and we're more than happy to help :)

OpenStudy (anonymous):

Why do they always leave when asked to show work :(

OpenStudy (anonymous):

It's just that it took me like three hours to complete all these answers and I already ripped up the scratch paper lol

OpenStudy (anonymous):

Would you write that on a test?

OpenStudy (anonymous):

Well the screenshot you took, lets go over that, how did you get c?

OpenStudy (anonymous):

It's alright if you don't understand a concept, math is difficult, and whenever you're ready to learn just tag us, until then adios!

OpenStudy (anonymous):

I just took a leap of faith and submitted the answers and I got most of them right so it's all good.

OpenStudy (triciaal):

very good you got most of them right but we want you to learn something not just get correct answers. there are some of us here that really want you to understand what to do. This will also make it easier for you when you get to higher math.